OpenStudy (anonymous):

doesn't 10 + 9 = 21 Really think about it

OpenStudy (igreen):

No, it doesn't. Really think about it. It still equals 19

OpenStudy (anonymous):

just think it pretty simple all it basically is is 9 is a number made up of 3x3 10 is a number made up of 5 x 2 were am i going witht this you will see then after you basically divide the number you get 3x3 and 5x2 multiply 3x5 =15 and then 3x2 =6 theres your answer 15+6 equals 21

OpenStudy (anonymous):

@iGreen :)

OpenStudy (igreen):

That makes no sense whatsoever.

OpenStudy (anonymous):

That makes no sense. -.-

OpenStudy (solomonzelman):

10 what? + 9 what? = ????? 10ft + 9inches 10meters + 9km on....

OpenStudy (solomonzelman):

what are the units?

OpenStudy (anonymous):

No matter how you cut, It equals 19. You can spin it around, wash it, throw it, and it'll still equal 19. Nice try though...

OpenStudy (anonymous):

its just numbers

OpenStudy (igreen):

That's like saying 10 is made out of 5 * 2, so 10 + 10 = 5 * 5 = 25

OpenStudy (igreen):

WHICH DOESN'T MAKE SENSE.

OpenStudy (anonymous):

XD i know i love screwing aruond with math :)

OpenStudy (briellshaw):

HE's talkiin bout the song 10+9=21

OpenStudy (anonymous):

i just got more people on here then i thought

OpenStudy (anonymous):

guys / girls i love you all

OpenStudy (anonymous):

i post this question so we can get along as family/friends related

OpenStudy (anonymous):

not to annoy you xD

OpenStudy (briellshaw):

https://m.youtube.com/watch?v=vsp18UOLveU here's the link to the song

OpenStudy (anonymous):

Post it more, it will be considered spam and you will get suspended. :)

OpenStudy (anonymous):

@briellshaw We don't care about any song, 9 + 10 doesn't equal 21. :)

OpenStudy (anonymous):

A song doesn't change the logic of math. :)

OpenStudy (anonymous):

@AnswerMyQuestions i changed it :)

OpenStudy (anonymous):

No you didn't. You just made up some mumbo jumbo.

OpenStudy (anonymous):

5 + 5 = 10 and 10 x 10 = 100 and 100 + 500 = 600, so 5 + 5 = 600. That's basically what you're saying.

OpenStudy (anonymous):

@AnswerMyQuestions smiling about someones suspensions and or wanting them is against rules thats being disrespectful :(

OpenStudy (anonymous):

Actually, I'm not smiling about anyone's suspension.

OpenStudy (anonymous):

and no i turned some stupid thing around to prove that it cant happen unless breaking rules of math :)

OpenStudy (anonymous):

Ok......

OpenStudy (anonymous):

We all know 9 + 10 doesn't equal 21, so you don't need to make a post of this. :)

OpenStudy (k_lynn):

10+9=19. I think that's pretty cut and dry. If I have 10 pencils and @AnswerMyQuestions gives me 9 more, I have 19 pencils.

OpenStudy (anonymous):

yes i know im not stupid

OpenStudy (anonymous):

I don't give away pencils, so you still have 10. jk LOL

OpenStudy (anonymous):

unlike you people who didnt read my post about i broke rules of math to solve so its a false equation

OpenStudy (anonymous):

and yes i called you idiots

OpenStudy (k_lynn):

Lol. So if I already have 10, and I go steal @AnswerMyQuestions 's 9 pencils, then I have 19

OpenStudy (anonymous):

u dont get it the post is over

OpenStudy (k_lynn):

Okie doke, see you :D